Please help !
if f(x)=1-x/1+x
then
f(-x)=? f(x+1)=? f(x)+1=? f(1/x)=? 1/f(x)=?
1
Expert's answer
2020-09-22T13:36:48-0400

Given : "\mathbf{f(x)=\dfrac{1-x}{1+x}}"


  1. "for\;f(-x)\; ,replace\;x\;by\;(-x)\;in\;f(x),\;then"


"\mathbf{f(-x)=\dfrac{1-(-x)}{1+(-x)}=\dfrac{1+x}{1-x}}"


2."\;for\;f(x+1)\;,replace\;x\;by\;(x+1)\;in\;f(x),\;then"


"\mathbf{f(x+1)=\dfrac{1-(x+1)}{1+(x+1)}=\dfrac{1-x-1}{1+x+1}=\dfrac{-x}{2+x}}"


3."\;for\;f(x)+1,\;we\;have-"


"\mathbf{f(x)+1=\dfrac{1-x}{1+x}+1=\dfrac{1-x+(1+x)}{1+x}=\dfrac{2}{1+x}}"


4."\;" "for\;f(1/x)\; ,replace\;x\;by\;(1/x)\;in\;f(x),\;then"


"\mathbf{f(1/x)=\dfrac{1-\dfrac{1}{x}}{1+\dfrac{1}{x}}=\dfrac{\dfrac{x-1}{x}}{\dfrac{x+1}{x}}=\dfrac{x-1}{x+1}}"


5."\;\;for\;1/f(x),\;we\;have-"


"\mathbf{\dfrac{1}{f(x)}=\dfrac{1}{\dfrac{1-x}{1+x}}=\dfrac{1+x}{1-x}}"